By using principle of mathematical induction prove 2^n > n^2

if n is an integer greater than 4.


1
Expert's answer
2021-09-24T00:59:12-0400

Let "P(n)" be the proposition that "2^n > n^2, n>4."

BASIS STEP: "P(5)" is true, because "2^5=32 > 25=5^2."


INDUCTIVE STEP: For the inductive hypothesis we assume that "P(k)" holds for an arbitrary integer "k>4." That is, we assume that


"2^k > k^2, k>4."

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"2^{k+1} > (k+1)^2"

We have that


"2^{k+1}=2(2^{k})>2k^2=k^2+k^2=k^2+k(k)"

">k^2+k(2+1)>k^2+2k+1=(k+1)^2, k>4"

This shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all integers "n>4."

That is, we have proven that "2^n > n^2" for all integers "n>4."